Oriole Company had $234,200 of net income in 2019 when the selling price per unit was $151, the variable costs per unit were $91, and the fixed costs were $575,800. Management expects per unit data and total fixed costs to remain the same in 2020. The president of Oriole Company is under pressure from stockholders to increase net income by $67,500 in 2020.

(a) Compute the number of units sold in 2019.

b) compute the number of units that would have to be sold in 2020 to reach the stockholders' desired profit level.

c) Assume that naylor company sells the same number of units in 2020 as it did in 2019. What would the selling price have to be in order to reacch the stockholders' desired profit level?
Answer #1

a.      Total contribution = Fixed cost + Net income

                                 = $575,800 + $234,200

                                 = $810,000

Contribution per unit = sales price – Variable cost

                                       = 151 – 91 = $60

Number of units sold = 810000 / 60 = 13500 units

b.      Required sales = (Fixed cost + Desired profit) /Contribution per unit

                           = (575800+234200+67500)/60

                            = 14625 units

c.       Sale price per unit = Total sales value/Number of sales units

                                 = (Variable cost+Fixed cost+Desired profit)/13500

                                  = (13500*91+575800+301700)/13500

                                  = $156.00